OpenStudy (anonymous):

5.Andrew claims that a wooden bookend in the shape of a 45°- 45°- 90° right triangle has side lengths of 9 in., 9 in., and 11 in. Is he correct? If so, show the work and if not, show why not.

OpenStudy (hba):

If it is a right angled triangle, then 9^2 + 9^2 should = 11^2 9^2 = 81 81 x 2 = 162 11^2 = 121

OpenStudy (hba):

162 is not equal to 121, therefore he is wrong

OpenStudy (anonymous):